# I-INR: Iterative Implicit Neural Representations

### v2

@@ 4 Experiments
To evaluate the effectiveness and robustness of our proposed method, we conduct extensive experiments across diverse tasks and backbones. Our evaluation encompasses regression tasks, such as 2D and 3D signal fitting, as well as generalization and robustness tasks, including image SR and image denoising. To demonstrate its comparative advantage, we benchmark our method against established baseline approaches, including SIREN [39], Gauss [30], and WIRE [36].
Implementation Settings: The proposed I-INR framework is implemented in PyTorch [25] and optimized using the Adam optimizer [10]. For all experiments, the noise parameter $\epsilon$ is empirically set to 0.1, and the inference $steps$ are fixed to 2 unless stated otherwise. The initial state $\mathcal{Z}$ is sampled from standard normal Gaussian distribution across all experiments. All baseline methods, including SIREN [39], WIRE [36], and Gauss [30], as well as their iterative counterparts, are initialized following the configurations specified in their respective papers. In I-INR, both the Auxiliary Network and the Projection layer employ lightweight MLPs to ensure computational efficiency while preserving expressivity.
Evaluation Metrics: To comprehensively evaluate our proposed method, we conduct extensive experiments using both distortion-based metrics, such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [45], as well as perceptual metrics like Learned Perceptual Image Patch Similarity (LPIPS) [48].
@@ 4.1 Results
Figure 5: Visual comparison of denoising results using various methods. The ground truth is compared against noisy input, baseline methods SIREN [39], WIRE [36], and Gauss [30], as well as their iterative counterparts. I-INR demonstrates superior artifact reduction and detail preservation compared to their non-iterative counterparts.
Table 1: Image fitting results on the Kodak dataset comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. The iterative approach outperforms baseline single-shot architectures by achieving higher reconstruction quality and enhanced fidelity.
SIREN [39] WIRE [36] Gauss [30]
@@
Image Fitting: For image fitting, we employ SIREN [39], WIRE [36], and Gauss [30], along with their iterative variants. Each network utilizes a three-layer MLP with 300 neurons per layer. In the case of iterative models, this corresponds to the Backbone INR illustrated in Figure 2(a). The networks are trained on the full-resolution Kodak dataset (768$\times$512 or 512$\times$768) [20], and the results, averaged across the entire dataset, are presented in Table 1. Our findings demonstrate that the proposed iterative models significantly outperform their one-shot counterparts. Figure 3 provides a qualitative comparison of all networks and their iterative adaptations on the kodim19 image. The error maps of baseline models exhibit homogeneous noise across the image, with more pronounced errors in high-frequency regions. In contrast, the iterative approaches yield smoother reconstructions with reduced errors throughout the image, including high-frequency areas. This ultimately results in a substantial improvement in fidelity, with PSNR gains of up to 2.7 dB observed for I-SIREN compared to its one-shot equivalent.
Image Super-resolution: We evaluate our I-INR framework against traditional one-shot INRs on SR tasks using 40 images from the DIV2K dataset [1]. The model is trained exclusively on 2×2$\times$2 × SR with a two-layer MLP of 256 neurons per layer. To assess generalization, we evaluate I-INR on both 2×2$\times$2 × and 4×4$\times$4 × SR, despite training only on 2×2$\times$2 ×. Table 2 compares its performance with single-shot baselines, demonstrating that our approach consistently enhances fidelity (measured as PSNR) and perceptual quality (measured as LPIPS) across all architectures. Figure 4 visualizes the butterfly image for 2×2$\times$2 × SR (please refer to supplementary material for visualization at 4×4$\times$4 × scale) , comparing each baseline with its iterative counterpart. The proposed approach consistently outperforms baseline methods, generating sharper reconstructions with fewer artifacts across all non-linearities. This improvement stems from the iterative process, which progressively refines the reconstruction, enabling more precise recovery of high-frequency details and allowing the model to generalize effectively beyond its training scale.
Table 2: Super-resolution (SR) results for 2×2$\times$2 × and 4×4$\times$4 × evaluations using a model trained exclusively on 2×2$\times$2 × SR. The table compares SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. PSNR (higher is better) and LPIPS (lower is better) are reported, with the best result in each column highlighted in bold.
Scale Method SIREN [39] WIRE [36] Gauss [30]
Factor PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
Table 3: Image denoising results on 40 images sampled from the DIV2K dataset [1], comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. The best values in each column are highlighted in bold.
SIREN [39] WIRE [36] Gauss [30]
@@
Figure 6: Visualization of 3D occupancy reconstruction results using various methods and their iterative counterparts. The ground truth is compared to baseline models (SIREN [39], WIRE [36], Gauss [30]) and their iterative counterparts. Iterative methods demonstrate superior reconstruction quality, achieving higher Intersection over Union (IOU) scores and capturing finer geometric details.
Image Denoising: To assess the robustness of I-INRs in modeling noisy signals, we utilize 40 high-resolution color images from the DIV2K dataset [1] and introduce Poisson-distributed noise to each pixel, with a maximum mean photon count of 30 and a readout noise level of 2, following  [36]. The same architecture and training procedure as SR is employed. Table 3 presents quantitative results comparing the denoising performance of SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. Across all three methods, the iterative models consistently achieve significant improvements in both PSNR and LPIPS over their single-shot baselines. Figure 5 provides a qualitative comparison of denoising performance for SIREN [39], WIRE [36], and Gauss [30] alongside their iterative versions. The iterative approach significantly enhances reconstruction quality across all architectures, achieving up to +3.25 dB PSNR improvement (for SIREN).
Refer to caption Refer to caption
(a) (b)
Figure 7: On the Kodak dataset for image fitting: (a) PSNR vs. training FLOPs comparison between SIREN [39] and I-SIREN. (b) PSNR vs. training iterations for I-SIREN with different learning rates.
Figure 8: Analysis of I-SIREN’s performance for super-resolution at 2$\times$ resolution across different iteration $steps$. (a) I-SIREN achieves peak PSNR at $steps=4$, where further increasing time steps enhances perceptual quality but reduces fidelity due to the fidelity-perception tradeoff. (b) Increasing the number of steps consistently improves perceptual quality. (c) PSNR reaches its highest performance at $steps=4$, and while additional iterations slightly reduce PSNR, it remains superior to the baseline.
3D Occupancy: The 3D occupancy task aims to reconstruct 3D shapes by modeling the occupancy of points in space. We conduct experiments on the Armadillo, Dragon, and Thai Statue 3D scenes [12], utilizing a two-layer MLP with 256 neurons per layer for both the baseline INRs and their iterative counterparts. Table 4 presents the averaged Intersection over Union (IoU) results, while Figure 6 visualizes the reconstructions achieved using different nonlinearities and their iterative variants. Notably, I-INRs outperform their single-shot baselines, generating sharper and more precise reconstructions. The iterative approach enhances detail preservation, particularly in complex regions. For instance, I-SIREN reconstructs significantly finer details at the base of the statue compared to the baseline SIREN, demonstrating its ability to model intricate spatial structures more effectively.
Table 4: Averaged Intersection over Union (IoU) results comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts on the Armadillo, Dragon, and Thai Statue 3D scenes.
SIREN[39$\uparrow$ WIRE[36$\uparrow$ Gauss[30$\uparrow$
@@ 4.2 Ablation Studies
Initial State: I-INR reconstructs a signal iteratively over multiple steps, starting from an initial state $\mathcal{Z}$. The choice of this initial state can significantly impact the reconstruction process. To assess its effect, we conduct extensive experiments with different types of initial states. Empirically, $\mathcal{Z}$ sampled from a standard normal Gaussian distribution $\mathcal{N}(0,1)$ consistently outperforms all-zeros and all-ones initial states across different non-linearities for the image fitting task. The results, summarized in Table 5, highlight the superiority of noise-based initial states in achieving higher PSNR and SSIM values.
Table 5: Impact of different initial states ($\mathcal{Z}$) on image fitting performance using the I-INR framework on the Kodak dataset. Noise-based initial states consistently achieve the best results across various non-linearities.
I-SIREN I-WIRE I-Gauss
Noise 37.53 0.961 33.73 0.924 31.93 0.884
Ones 37.42 0.961 33.33 0.921 30.96 0.867
Zeros 37.47 0.961 33.39 0.921 31.22 0.882
Training and Inference Complexity: Our proposed I-INR method achieves faster convergence than the non-iterative baseline while maintaining similar training complexity, as shown in Figure 7a. The results are averaged over the Kodak dataset for the image fitting task. Figure 7b further illustrates I-SIREN’s training performance under different learning rates.
While I-SIREN converges faster, this improvement comes at a slightly higher per-iteration computational cost, requiring 107.2 GFLOPS compared to 106.8 GFLOPS for SIREN. Additionally, since I-SIREN reconstructs images iteratively, the total inference cost scales with the number of inference steps. In our experiments, we consistently use $steps=2$, resulting in a total inference complexity of 214.4 GFLOPS for I-SIREN, compared to 106.8 GFLOPS for SIREN.
Number of Layers: For the image fitting task on the Kodak dataset, we investigate the impact of increasing the number of layers in SIREN and its iterative counterpart, I-SIREN. The results, presented in Table 6, show that while the performance of SIREN saturates beyond a certain depth, specifically at five layers, I-SIREN continues to improve as more layers are added. Notably, a three-layer I-SIREN outperforms a five-layer SIREN, demonstrating the efficiency of the iterative approach.
Table 6: Impact of varying layer depths on the image fitting task using SIREN [39] and I-SIREN on the Kodak dataset. I-SIREN consistently outperforms SIREN [39], with a three-layer I-SIREN achieving higher PSNR than a five-layer SIREN [39].
Layers I-SIREN SIREN
PSNR Params (k) PSNR Params (k)
3 37.53 274 34.62 273
4 38.71 364 36.60 363
5 39.01 455 36.49 453
Analysis of Reconstruction Steps: We analyze the impact of reconstruction steps on 2$\times$ super-resolution using 40 images sampled from the DIV2K dataset, as shown in Figures 8b and 8c. Our findings reveal that the model achieves its highest PSNR at $steps=4$, with the most significant gain occurring when $steps$ increases from 1 to 2. Beyond $steps=2$, the PSNR begins to plateau. Notably, although PSNR decreases for $steps>4$, the perceptual quality (measured by LPIPS) continues to improve. This behavior aligns with the Perception-Distortion Tradeoff [2], where introducing additional high-frequency details enhances realism but slightly deviates from the ground-truth, thus lowering distortion-based metrics such as PSNR. Figure 8a presents a visual comparison of SIREN and I-SIREN on an image from DIV2K dataset at various time steps. I-SIREN reconstructions become progressively more vibrant and perceptually appealing as step increases, confirming that additional refinement steps are beneficial for capturing finer textures and high-frequency components. This trend holds for other reconstruction tasks as well, as detailed in our supplementary material.
@@ 6.1 Additional Quantative Results
Table 7 presents a detailed comparison of our proposed method against the baseline across image fitting, denoising, and SR at 2×2$\times$2 × for all three evaluation metrics (PSNR, SSIM, and LPIPS). Our iterative approach consistently enhances performance across tasks and different baselines, achieving superior reconstruction quality.
@@ 6.2 Additional Qualitative Comparison
We also present an additional qualitative comparison of our proposed method against SIREN, Gauss, and WIRE. Figures 9101112, illustrates a comparative analysis for image fitting, SR at 2×2$\times$2 × and 4×4$\times$4 × scales, and image denoising, demonstrating the improvements achieved by our approach over baseline INRs. Our proposed method consistently outperforms single-shot baselines, demonstrating superior reconstruction quality across various models and tasks.
@@ 7 Additional Ablation Studies
Additional Reconstruction Steps Analysis: For SR at 2×2$\times$2 ×, we further analyze the reconstruction quality of I-SIREN over multiple steps, comparing it to SIREN and the ground truth. As shown in Figure 13, the results demonstrate that with an increasing number of steps, details become more pronounced, and more high-frequency information is incorporated into the reconstruction.
@@
We also analyze the effect of reconstruction steps on image denoising performance using 40 images from the DIV2K dataset (see Figure 15). Our findings indicate that PSNR peaks at $steps=2$ before gradually declining as $steps$ increases. Meanwhile, LPIPS improves until $step=4$ and stabilizes around step 5. Beyond $steps=4$, both PSNR and LPIPS degrade as I-INR begins reconstructing noise present in the data. At higher iterations, the model inadvertently reconstructs these distortions, leading to a decline in performance.
Feature Fusion: We conduct ablation studies on the final feature-fusion mechanism for signal reconstruction, evaluating two strategies that combine the outputs of the Backbone and Projection modules (Figure 2a of main manuscript):
[…]
In Multiplicative fusion, the final output $\mathcal{I}(x)_{t}$ is obtained via element-wise multiplication of the Backbone and Projection outputs. The Adaptive method, conversely, uses the interpolation factor $t$ to smoothly shift emphasis from the Backbone output (dominant at $t=1$) toward the Projection output at higher values of $t$. Quantitative results comparing both fusion strategies for the image-fitting task are summarized in Table 8. Multiplicative fusion consistently outperforms Adaptive fusion, as evidenced by the results. Therefore, we adopt Multiplicative fusion as the standard approach for our method.

### v3

@@ 5 Experiments
To evaluate the effectiveness and robustness of our proposed method, we conduct extensive experiments across diverse tasks and backbones. Our evaluation encompasses regression tasks, such as 2D and 3D signal fitting, as well as generalization and robustness tasks, including image SR and image denoising. To demonstrate its comparative advantage, we benchmark our method against established baseline approaches, including SIREN, Gauss, and WIRE.
Implementation Settings.
The proposed I-INR framework is implemented in PyTorch (paszke2019pytorch) and optimized using the Adam optimizer (diederik2014adam). For all experiments, the noise parameter $\epsilon$ is empirically set to 0.1, and the inference stepssteps are fixed to 2 unless stated otherwise. The initial state $\mathcal{Z}$ is sampled from a standard normal Gaussian distribution across all experiments. All baseline methods, including SIREN, WIRE, and Gauss, as well as their iterative counterparts, are initialized following the configurations specified in their respective papers. In our implementation, both FeedbackNet and FuseNet are lightweight MLPs with two layers, each layer having a width of 30 for FeedbackNet and 100 for FuseNet.
Evaluation Metrics.
To comprehensively evaluate our proposed method, we conduct extensive experiments using both distortion-based metrics, such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) (wang2004image), as well as perceptual metrics like Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable). For a detailed analysis, refer to the supplementary materials.
@@ 5.1 Results
SIREN WIRE Gauss
@@
Table 1: Image fitting results on Kodak comparing SIREN, WIRE, and Gauss with their iterative counterparts.
Image Fitting.
For image fitting, each baseline employs a 3-layer MLP with 300 neurons per layer, serving as the Backbone INR for iterative models (Figure 2(a)). All networks are trained on the full-resolution Kodak dataset (kodak), with average results reported in Table 1. The proposed iterative models consistently outperform their one-shot counterparts. As shown in Figure 3, error maps from baseline models display uniform noise, especially in high-frequency regions, whereas iterative variants produce smoother reconstructions with reduced errors, including in challenging areas.
Image Super-resolution.
We evaluate our I-INR framework against traditional one-shot INRs on SR tasks using 40 images from the DIV2K dataset (agustsson2017ntire). The model is trained exclusively on $2\times$ SR with a two-layer MLP of 256 neurons per layer. To assess generalization, we evaluate I-INR on both $2\times$ and $4\times$ SR, despite training only on $2\times$. Table 2 compares its performance with single-shot baselines, demonstrating that our approach consistently enhances fidelity (PSNR) and perceptual quality (LPIPS) across all architectures. Figure 4 visualizes the butterfly image for $2\times$ SR (refer to supplementary for visualization at $4\times$), comparing each baseline with its iterative counterpart. The proposed approach consistently outperforms baseline methods, generating sharper reconstructions with fewer artifacts across all non-linearities.
Scale Method SIREN WIRE Gauss
@@
Table 2: Comparison of (SR) performance for $2\times$ and $4\times$ upscaling using a model trained on $2\times$ SR. Results are shown alongside SIREN, WIRE, and Gauss, including their iterative counterparts. The best results are shown in bold.
SIREN WIRE Gauss
@@
Table 3: Image denoising results on 40 images sampled from the DIV2K dataset, comparing SIREN, WIRE, and Gauss with their iterative counterparts.
Figure 6: Visualization of 3D occupancy reconstruction results using various methods and their iterative counterparts. The ground truth is compared to baseline models SIREN, WIRE, Gauss, and their iterative counterparts.
Image Denoising.
To assess the robustness of I-INRs in modeling noisy signals, we utilize 40 high-resolution color images from the DIV2K dataset and introduce Poisson-distributed noise to each pixel, with a maximum mean photon count of 30 and a readout noise level of 2, following  (saragadam2023wire). The same architecture and training procedure as SR is employed. Table 3 presents quantitative results comparing the denoising performance of baseline architectures with their iterative counterparts. Iterative models consistently achieve significant improvements in both PSNR and LPIPS over their single-shot baselines. Figure 5 provides a qualitative comparison of denoising performance for SIREN, WIRE, and Gauss alongside their iterative versions. The iterative approach significantly enhances reconstruction quality across all architectures, achieving up to +3.25 dB PSNR improvement (for SIREN).
[…]
3D Occupancy.
The 3D occupancy task aims to reconstruct 3D shapes by modeling the occupancy of points in space. We conduct experiments on the Armadillo, Dragon, and Thai Statue 3D scenes (gal2006salient), utilizing a two-layer MLP with 256 neurons per layer for both the baseline INRs and their iterative counterparts. Table 4 presents the averaged Intersection over Union (IoU) results, while Figure 6 visualizes the reconstructions achieved using different nonlinearities and their iterative variants. Notably, I-INRs outperform their single-shot baselines, generating sharper and more precise reconstructions. The iterative approach enhances detail preservation, particularly in complex regions as shown in Figure 6.
SIREN $\uparrow$ WIRE $\uparrow$ Gauss $\uparrow$
@@
Table 4: Averaged Intersection over Union (IoU) results comparing SIREN, WIRE, and Gauss with their iterative counterparts on the Armadillo, Dragon, and Thai Statue 3D.
@@ 5.2 Ablation Studies
We ablate inference complexity, Feedback/FuseNet impact, and reconstruction steps. Further analysis on initial state, training cost, layer depth, and statistical significance is included in the supplementary material.
Inference Complexity.
During reconstruction, I-SIREN executes the Backbone network only once, while subsequent refinement steps involve forward passes through the lightweight FeedbackNet and FuseNet modules. The iterative setup introduces minimal computational overhead: the Backbone accounts for 106.8 GFLOPs, and each refinement step adds just 0.43 GFLOPs. For two refinement steps ($\text{steps}=2$), the total computational cost amounts to 107.6 GFLOPs, approximately equivalent to the original SIREN.
Impact of FeedbackNet and FuseNet:
To evaluate the impact of architectural components of I-INR, we performed ablation studies by selectively disabling the FeedbackNet and FuseNet modules and varying their parameter scales. These experiments were conducted on the image fitting task using I-SIREN on the Kodak dataset. As summarized in Table 5, removing both modules leads to a significant performance drop, with PSNR falling to 20.27. Introducing FeedbackNet alone substantially improves reconstruction quality, and the addition of FuseNet provides further gains. The best performance is achieved when both modules are included, with moderate parameter scaling yielding an optimal balance between computational cost and accuracy.
FeedbackNet (Params) FuseNet (Params) PSNR
20.27
[…]
Table 5: Impact of FeedbackNet and FuseNet on image fitting task with I-SIREN on the Kodak dataset.
Analysis of Reconstruction Steps.
We evaluate the effect of reconstruction steps on $2{\times}$ super-resolution using 40 images from the DIV2K dataset (Figures 7b and 7c). PSNR peaks at $steps=4$, with the largest improvement from 1 to 2 steps; beyond 2, gains saturate. While PSNR slightly drops for $steps>4$, perceptual quality (LPIPS) continues to improve, reflecting the Perception-Distortion Tradeoff (blau2018perception). Figure 7a visually compares SIREN and I-SIREN across steps, showing that I-SIREN produces increasingly vivid and detailed reconstructions, confirming that additional refinement steps are beneficial for capturing finer textures and high-frequency components. This trend generalizes to other tasks, as discussed in the supplementary material.
@@ 7.1 Additional Quantative Results
Table 6 presents a detailed comparison of our proposed method against the baseline across image fitting, denoising, and SR at $2\times$ for all three evaluation metrics (PSNR, SSIM, and LPIPS). Our iterative approach consistently enhances performance across tasks and different baselines, achieving superior reconstruction quality.
@@ 7.2 Additional Qualitative Comparison
We also present an additional qualitative comparison of our proposed method against SIREN, Gauss, and WIRE. Figures 891011, illustrates a comparative analysis for image fitting, SR at $2\times$ and $4\times$ scales, and image denoising, demonstrating the improvements achieved by our approach over baseline INRs. Our proposed method consistently outperforms single-shot baselines, demonstrating superior reconstruction quality across various models and tasks.
@@ 8 Additional Ablation Studies
Initial State.
I-INR reconstructs a signal iteratively over multiple steps, starting from an initial state $\mathcal{Z}$. The choice of this initial state can significantly impact the reconstruction process. To assess its effect, we conduct extensive experiments with different types of initial states. Empirically, $\mathcal{Z}$ sampled from a standard normal Gaussian distribution $\mathcal{N}(0,1)$ consistently outperforms all-zeros and all-ones initial states across different non-linearities for the image fitting task. The results, summarized in Table 7, highlight the superiority of noise-based initial states in achieving higher PSNR and SSIM values.
Training Complexity. Our proposed I-INR method achieves faster convergence than the non-iterative baseline while maintaining similar training complexity, as shown in Figure 12a. The results are averaged over the Kodak dataset for the image fitting task. Figure 12b further illustrates I-SIREN’s training performance under different learning rates.
Number of Layers in Backbone Network. For the image fitting task on the Kodak dataset, we investigate the impact of increasing the number of layers in SIREN and its iterative counterpart, I-SIREN. The results, presented in Table 8, show that while the performance of SIREN saturates beyond a certain depth, specifically at five layers, I-SIREN continues to improve as more layers are added. Notably, a three-layer I-SIREN outperforms a five-layer SIREN, demonstrating the efficiency of the iterative approach.
Additional Reconstruction Steps Analysis. For SR at $2\times$, we further analyze the reconstruction quality of I-SIREN over multiple steps, comparing it to SIREN and the ground truth. As shown in Figure 13, the results demonstrate that with an increasing number of steps, details become more pronounced, and more high-frequency information is incorporated into the reconstruction.
@@
We also analyze the effect of reconstruction steps on image denoising performance using 40 images from the DIV2K dataset (see Figure 15). Our findings indicate that PSNR peaks at $steps=2$ before gradually declining as stepssteps increases. Meanwhile, LPIPS improves until $step=4$ and stabilizes around step 5. Beyond $steps=4$, both PSNR and LPIPS degrade as I-INR begins reconstructing noise present in the data. At higher iterations, the model inadvertently reconstructs these distortions, leading to a decline in performance.
Feature Fusion. We conduct ablation studies on the final feature-fusion mechanism for signal reconstruction, evaluating two strategies that combine the outputs of the Backbone and FuseNet modules (Figure 2a of main manuscript):
$\mathbf{b}:=\operatorname{Backbone}(x),\qquad\mathbf{f}:=\operatorname{FeedbackNet}(\hat{g}(x)_{t},\,x,\,t)$
$\mathbf{z}:=\operatorname{FuseNet}(\operatorname{concat}(\mathbf{f},\,\mathbf{b}))$
$\text{Multiplicative:}\quad\mathit{out}=\mathbf{b}\odot\mathbf{z},$
$\text{Adaptive:}\quad\mathit{out}=\mathbf{b}\cdot t+\mathbf{z}\cdot(1-t)$
In Multiplicative fusion, the final output is obtained via element-wise multiplication of the Backbone and FuseNet outputs. The Adaptive method, conversely, uses the interpolation factor $t$ to smoothly shift emphasis from the Backbone output (dominant at $t=1$) toward the FuseNet output at higher values of $t$. Quantitative results comparing both fusion strategies for the image-fitting task are summarized in Table 9. Multiplicative fusion consistently outperforms Adaptive fusion, as evidenced by the results. Therefore, we adopt Multiplicative fusion as the standard approach for our method.
Time Complexity. We compare SIREN and I-SIREN on the Kodak dataset, trained for 2000 iterations on an NVIDIA RTX 4090 GPU. As shown in Table 10, I-SIREN adds only a small overhead in training time ($6\%$) and inference latency (+3+3 ms per image) while achieving higher reconstruction quality, confirming its efficiency.
Statistical Significance. All prior experiments were conducted using a single random seed. To assess the robustness and statistical significance of our approach, we additionally evaluate performance across five different seeds. Table 11 reports the results for image fitting, SR, and denoising tasks, comparing SIREN and I-SIREN, along with confidence intervals. The dataset and experimental settings remain consistent with those previously described. As shown in Table 11, I-SIREN consistently outperforms SIREN across all tasks. The improvements achieved by I-SIREN are statistically significant, with p-values << 0.00001 based on the Wilcoxon signed-rank test.
@@
I-SIREN I-WIRE I-Gauss
Noise 37.53 0.961 33.73 0.924 31.93 0.884
Ones 37.42 0.961 33.33 0.921 30.96 0.867
Zeros 37.47 0.961 33.39 0.921 31.22 0.882
Table 7: Impact of different initial states ($\mathcal{Z}$) on image fitting task using the I-INR framework on Kodak dataset.
Layers I-SIREN SIREN
PSNR Params (k) PSNR Params (k)
3 37.53 274 34.62 273
4 38.71 364 36.60 363
5 39.01 455 36.49 453
Table 8: Impact of varying Backbone network layer depths on the image fitting task using SIREN and I-SIREN on the Kodak dataset.
@@
Method Train Time (s) Inference (ms / image) PSNR (dB)
SIREN 100 18.9 34.57
I-SIREN (2 steps) 106 21.9 37.53
Table 10: Training and inference time comparison on the Kodak dataset (RTX 4090).
Method Image Fitting SR Denoising
I-SIREN 37.50 $\pm$ 0.247 27.22 $\pm$ 0.045 25.36 $\pm$ 0.044
SIREN 34.59 $\pm$ 0.027 26.77 $\pm$ 0.035 23.89 $\pm$ 0.021
Table 11: Evaluation on 5 different seeds for image fitting, super-resolution tasks, and denoising on SIREN and I-SIREN.